OpenStudy (anonymous):

solve by the square root method (x-3)^2+33=0 not sure if my answer is correct I got x=+-the square root of 6times 7

OpenStudy (inkyvoyd):

The answershould be a complex number.

OpenStudy (cwtan):

what is meant by square root method?

OpenStudy (anonymous):

This has no real roots. It has 2 complex roots \[(x-3)^2 =-33 = i^2 33\\ x-3 =\pm i \sqrt {33}\\ x = 3\pm i \sqrt {33}\\ \]